- Press Release
- Nov 26, 2022
NASA Space Station On-Orbit Status 1 Apr 2004
All ISS systems continue to function nominally, except those noted previously or below. (April Fool’s Day message to the crew: “We’re going to have you go EVA tomorrow for some pictures of the TRRJ…”)
CDR/SO Michael Foale completed the regular periodic inspection of smoke detectors in the Lab module, Airlock (A/L) and Node.
Later, Mike also conducted the monthly PEP (portable emergency provisions) inspection, verifying that PFEs (portable fire extinguishers), PBAs (portable breathing apparatus), QDMs (quick-don masks) and the oxygen extension hose tee kit are all free of damage, to ensure their functionality. [There are a total of five PBAs in the USOS, viz., two in Node lockers, two in Lab module forward and aft endcone lockers, and one in the Airlock PEP locker.]
FE Alexander Kaleri performed the periodic leakage inspection of the two foam applicators in the ISS Leak Patch Kit. [The inspection is done wearing protective gear in case the foam substance, of toxicity level 1, has leaked from the applicator inside the kit.]
The crew began equipment transfer to and loading operations in Progress-260/13P for disposal, using a newly uplinked list of hardware identifications, ISS stowage locations, and recommended accommodations in the cargo ship-turned-trash remover.
Kaleri completed requested photography of window #13 in the Service Module (SM), which on 3/19 was found by the crew to have acquired a new impact crater of about 2.5 mm diameter and a scratch. [The defects were not present on previous pictures. Alex took several shots, with a ruler taped to the window frame for scale reference. Window #13 is one of three in the small conical transition segment from the PkhO Transfer Compartment, measures 22.8 cm in diameter and is protected by a manually driven external cover.]
After Mike Foale had reconnected the UOP-DCP (utility outlet panel-to-display & control panel) bypass power cable at the Lab RWS (robotics workstation) in the morning and the ground had powered up the MSS (Mobile Service System) and its video equipment, Foale took the SSRMS (Space Station Remote Manipulator System) through the scheduled proficiency training session. One hour was reserved for set up and review of the Dynamic Operational Ubiquitous Graphics (DOUG) application and for general preparations, leaving two hours for the actual operations, which were executed without hitch. [The procedure’s objective was an outside inspection with the MSS video cameras of the CDRA/MISSE area, to check for ice collection around the CDRA (CO2 Removal Assembly) vent nozzle and inspect the MISSE (Materials ISS Experiment), extensive surveys of the ISS nadir side, inspection of more H-Fixtures on the P1 truss outboard and inboard, and inspection of the P6 forward PVR (photovoltaic radiator), where a bubble has been discovered under the fabric on the base structure. The surveys were recorded on board for downlink later. MCC-H managed the VTRs (video tape recorders) and routed to them.]
Alex Kaleri ran a test on the new LT1 (laptop 1) at the SM Central Post, after initializing the laptop’s BIOS settings, to check on its network connectivity, i.e., its communication with the Russian KTsP1 (Central Post Computer 1). [Depending on the outcome, Sasha was to either turn LT1 off as per nominal deactivation procedure, or to replace it with LT3 and activate the latter. Early indications are that the test went “nominal”.]
Later, the FE performed the regular once-a-week maintenance reboot on the operational portable (PCS) laptops and also restarted the SSC OCA comm router laptop (normally every two weeks).
Sasha downloaded data and imagery collected of the BIO-5 Rasteniya-2 (“Plants-2”) experiment to the computer for subsequent downlink to the ground. [Rasteniya studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.]
Sasha also attended to the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities).
At 3:55am EST, the crew downlinked greetings and answers to questions for replay on the “Good Morning” news (vesti) program on the Rossia TV channel in Moscow. [“Six months spent on the ISS are behind you. Which month was the most challenging, and which was the most successful and most interesting to you?”; “What do you think about space tourism? Do you need to have a dedicated crewmember as a full-time ISS tour guide?” and finally: “Spasibo! Do svidaniya!”]
At 12:25pm, the crew conducted an interactive educational PAO exchange with students at the Center of Science and Industry (COSI) in Columbus, OH, which is headed by former Shuttle astronaut (and first US woman to conduct an EVA) Dr. Kathryn Sullivan. [Student questions had been uplinked beforehand, and the crew had prepared tools and a tools block to demonstrate their use. (“How much force is needed to drive a nail into a piece of wood and how does it differ from a hammer used on Earth?”; “How is using a screwdriver (or a wrench) in space different than on Earth?”]
Later in the day, at 1:25pm, CDR Foale had a 10-min. amateur/ham radio exchange with pupils at two schools, Jacques Prevert and Georges Brassens, in Saint Mard, France. [The schools have set up a technical/scientific workshop called “Syntheses 3D” last year to allow the students to get “hands on” experience with satellites and the ISS, by learning radio technology and using computers equipped with 3D animation and modeling software.]
The CDR conducted his periodic teleconference with the Increment 9 crew (Gennady Padalka, Mike Fincke, Andrë Kuipers), via audio/phone patch on S-band.
Mike and Sasha conducted their regular daily physical exercise program of 2.5 hrs on TVIS treadmill, CEVIS bike, RED exerciser and VELO cycle with load trainer. Kaleri received special recommendations from TsUP on the use of the TVIS with SLDs (subject load devices) and the Russian TNK loading suit, with the desired axial load setting during the final flight stage.
At 1:30pm, MCC-Houston remote-commanded a swap ofthe U.S. guidance, navigation & control computers (GNC MDMs) from primary to backup, to upload a software patch to the GNCs that will allow them to accept more RGA (rate gyro assembly) rate data, thus improving the overall performance of the US attitude determination system. [Both MDMs were loaded, then reconfigured and recovered. Preparatory to the swap-out, at 12:55pm, attitude control authority was handed over to the Russian MCS (motion control system), which maneuvered the station to LVLH minimum prop attitude, then to LVLH TEA (torque equilibrium attitude), prior to returning authority directly to the U.S. momentum management controllers at 4:20pm, using the new CMG-protecting procedure.]
TsUP/Moscow is conducting a test of the Regul-Paket comm system, to troubleshoot problems of the “loopback mode”, i.e., the return link from ISS to TsUP. Crew involvement was/is not required, except for Kaleri starting up the Paket laptop with the appropriate application this morning. Rebooting the laptop tomorrow morning will automatically restore normal ops.
U.S. testing of the TRRJ (thermal radiator rotary joint) software continues, and lessons are being learned. [Specialists are currently looking into the repeated failure of the software to switch over from thermal loop A to loop B when the responsible FDIR (failure detection, isolation, and recovery) system perceives a string failure (in this case test-commanded, i.e., simulated).]
The BCAT-3 (Binary Colloidal Alloy Test 3) experiment has progressed at a much faster rate of the phase separation of the samples than anticipated, and the investigators are hoping for an additional photo session to better understand the time scale at work within them.
Correction to yesterday’s report on the Orlan-M checkout: Increment 9 will feature only two (not three) Russian Orlan EVAs, to wit:
- EVA-10 (7/22/04) — for installing protective elements on the DC-1’s circular handrails as well as soft handrails on DC-1, conducting experiment activities (Kromka, Platan-M), and preparing for ATV docking (part 1);
- EVA-11 (8/24/04) — for replacing a removable panel for a liquid flow regulator on the FGB, installing four fairleads on FGB handrails as well as a plume impingement & deposit monitoring unit, and preparing for ATV docking (part 2).
Today’s CEO (Crew Earth Observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Ganges River Basin (looking eastward, to the right of track, for oblique views of the regional smog and haze layer. Trying for views which indicate the thickness of these aerosols as referenced against the darker wall of the Himalayan Massif), Amman, Jordan (a nadir pass over the Jordanian capital located about 40 miles northeast of the Dead Sea), Brasilia, Brazil (the pass should have been early enough to beat the cumulus field produced by daytime heating. Looking just right of track for the unusual pattern of this city on a lake), and Chicago, Illinois (the Windy City forecast was for clearing by the time of the ISS pass. Looking just right of track as the station approached Lake Michigan from the SW.)
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at
U.S. and Russian Segment Status (as of today, 2:51pm EST).
- Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (repair not completed; to be tested ASAP — see above note).
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.5; ppO2 (mmHg) — 159.4; ppCO2 (mmHg) — 3.7;
- SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C) — 19.8.
- FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 22.7.
- Node: Pressure (mmHg) — 752.79; temperature (deg C) — 22.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 754.92; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 755.02; temperature (deg C) — 22.4; shell heater temp (deg C) — 22.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- (n/a = data not available)
- PMA-1: Shell heater temp (deg C) — 22.3
- PMA-2: Shell heater temp (deg C) — 19.9
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind mode, non solar-tracking, drag reduction-biased).
- SM batteries: Battery #8 is off-line; all other batteries (7) are in “Partial Charge” mode.
- FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); battery #5 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- &C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup.
- INT-2 is operating; INT-1 is Off.
- EXT-1 is On (primary), EXT-2 is Off.
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
- FGB MDM-1 is Off (failed); MDM-2 is Operational.
- Total propellant load available: 3926 kg (8655 lb) as of 3/26/04; [SM(755) + FGB(2512) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 3 CMGs on-line (CMG-1 failed).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
- XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- S-band is operating nominally (on string 2).
- Ku-band is operating nominally.
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
- MBS: KA (keep alive) power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 7:00am EST [= epoch]):
- Mean altitude — 365.6 km
- Apogee — 372.8km
- Perigee — 358.3 km
- Period — 91.86 min.
- Inclination (to Equator) — 51.63 deg
- Eccentricity — 0.0010715
- Orbits per 24-hr. day — 15.68
- Mean altitude loss last 24 hours — 155 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 30628
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see